OpenStudy (anonymous):

A given amount of water cools from 68.4C to 56.7C. During this cooling process, 12.4 kcal are released into the surrounding area. What is the water’s mass? will medal

OpenStudy (aaronq):

\(q=m*C_p*\Delta T\) Solve for m

OpenStudy (anonymous):

m is mass, delta t is change in temperature, what is q and Cp?

OpenStudy (aaronq):

q is the heat (thermal energy) and \(C_P\) is the specific heat capacity

OpenStudy (anonymous):

12.4 = m * 4.179 * 11.7 m = .2536 does this look right?

OpenStudy (anonymous):

.2536 grams or kilograms?

OpenStudy (aaronq):

what are the units on the heat capacity?

OpenStudy (aaronq):

they were probably "\(J/g ^oC\)" so before you use 12.4 kcal, you need to convert them to joules. 1 kcal = 4184 J

OpenStudy (anonymous):

51,881.6 = m * 4.179 * 11.7 51,881.6 = m * 48.8943 mass = 1,061.1 g does this look better?

OpenStudy (aaronq):

yep!

OpenStudy (anonymous):

cool thanks for your help

OpenStudy (aaronq):

no problem